OpenStudy (anonymous):

1.6 as a fraction. Please help! @texaschic101

OpenStudy (anonymous):

1 and 3/5 or 5/8?

OpenStudy (anonymous):

well .6 as a fraction is 6/10 or it can be simplified to 3/5. so 1 and 3/5 or 1 and 6/10

OpenStudy (anonymous):

medal and fan

OpenStudy (texaschic101):

look at what place value the 6 is in. It is in the tenths place...your 1 is on the left side of the decimal, so it is a whole number. 1 6/10 = 16/10 which reduces to 8/5 or 1 3/5

OpenStudy (anonymous):

8/5

OpenStudy (anonymous):

but that is not part of the answers

OpenStudy (texaschic101):

1 6/10 = 1 3/5 = 8/5...they all equal

OpenStudy (anonymous):

so then which one is it?

OpenStudy (texaschic101):

do you have answer choices ? Because 8/5 is an improper fraction and 1 3/5 is a mixed number fraction

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

The Martinez family walked three over five of a mile in three over eight of an hour. What equation can be used to calculate their unit rate in miles per hour? three-fifths over three-eights equals one and three-fifths miles per hour three-eighths over three-fifths equals one and three-fifths miles per hour three-eighths over three-fifths equals five over eight mile per hour three-fifths over three-eights equals five over eight mile per hour

OpenStudy (anonymous):

answer?

OpenStudy (texaschic101):

(3/5) / (3/8) = 3/5 * 8/3 = 24/15 miles per hr which is also 1 9/15 or 1 3/5 miles per hr....its A

OpenStudy (anonymous):

ok thanks so much! as i said im your fan and i gave you a medal! :D

OpenStudy (anonymous):

i gave you pretty much the answer from the beginning...